Lens before pre-cavity BN detector is PLCX- 25.4-25.6-C-1064 (f = 50 mm), this wasn't quite at the right distance. I removed lens and PD (Newport 1611) and looked at the overlap at large distance and fine tuned the MM. After reinstalling PD and lens I got a DC voltage of -2.8 V (equivalent to 200 µW of power when I account for 1 MΩ load of oscilloscope). So maybe position of PD in the focus could be better still.
I see a beat note of -80 dBm on a SA noise floor of ~100 dBm: enough for diagnostics now. This detector would perform much better with proper mode matching of the two paths. But science return on investment isn't worth it now.
We get a beat note of ~62 MHz for
South slow controls voltage = 0.8052 V
North slow controls voltage = 3.9250 V
Currently The north is locking at 3.92 V and the south at around 0.66. So some fine tuning of the north cavity temperature is needed to produce a BN in the 125 MHz band width of the 1811 detector on the transmission of the cavities.
Laser temperature tuning goes 4 GHz/V applied at the laser head.
The current BN with both caivities locked on thier closest resonances is 456 MHz (south slow = 0.6597 V, north slow = 3.9109 V). I turned north heater voltage off from 11.5 V to 0 V to test which direction the heating needed to go.
This was the wrong direction. Going back the other way turned voltage back up to 12.0 V. This will take a while